# Weak Convexity of Fisher Information Matrix and Superresolved Localization of Blinking Sources of Light

### v2

@@ I Introduction
$\left(\Delta x\right)^{2}\geq 1/\bm{F},$ (1)
where $\bm{F}$ is the Fisher information on this parameter contained in the measurement data. Here and below we use bold letters for denoting quantities related to the entire measurement session typically consisting of many photodetections. In the case of localization of a single emitter, the subsequent photodetections are statistically independent, and the total Fisher information carried by $N$ photons is simply $\bm{F}=NF_{\text{1phot}}$, where $F_{\text{1phot}}$ is the Fisher information on the emitter position carried by one photon. The Cramér-Rao bound can be asymptotically reached in the high $N$ limit when the estimator is given by the maximum likelihood estimation [18, 19]. Thus, the $1/\sqrt{N}$ scaling of the position standard deviation is a mere consequence of the additivity of Fisher information with respect to independent observations. The single-photon Fisher information can be calculated for various models of the PSF and the accompanying noise [20, 21] providing the Cramér-Rao bound for any given $N$. The variance of the estimated emitter position was shown to approach the Cramér-Rao bound for $N=10^{2}-10^{4}$ photons allowing thus for superresolution.
The same argument can be applied to determining the separation of two closely spaced emitters: the total Fisher information carried by $N$ photons on the separation is $\bm{F}=NF_{\text{sep,1phot}}$, where $F_{\text{sep,1phot}}$ is the information on the separation carried by one photon. This information tends to zero together with the separation of the emitters [22], however, for whatever small $F_{\text{sep,1phot}}$, one can make the total information arbitrarily large by collecting sufficiently large number of photons and thus determine the separation with arbitrary precision. In practice, a fluorophore emits a limited number of photons before photobleaching, and, therefore, individual localization of blinking emitters gives much better results than a simultaneous localization of multiple cofluorescent emitters within the width of PSF. The precision gain attained by making the sources blink is the key feature of SMLM. A question arises: Is it a general law of Nature that blinking emitters send more information on their locations than cofluorescent ones?
@@
The paper is structured as follows. In Sec. II we recast the formalism of quantum metrology [23, 24, 25, 26] in the context of localization of dim incoherent sources of light [22, 27, 28, 29, 30], extending a more traditional approach of classical metrology typically used in this context [20, 21, 31, 32, 33, 34], and introduce three measures of localization efficiency. In Sec. III, we prove the fundamental properties of the Fisher information matrix mentioned above. In Sec. IV, we compare two scenarios of cofluorescent and blinking sources and show that the latter always has an informational advantage. We illustrate the developed theory by an example of two incoherent sources in Sec. V and extend the results to the quantum Fisher information matrix in Sec. VI. The obtained results are summarized in Conclusions.
@@ II.1 Cramér-Rao bound
The objective of quantum metrology is inference of an unknown parameter $\theta$, which may be a vector $\theta=\{\theta_{1},...,\theta_{M}\}$, encoded into the state of the object $\rho_{O}(\theta)$, which is a density operator in the Hilbert space of the object. For this purpose a generalized quantum measurement of the object is repeated $N$ times, the object being prepared in the same state $\rho_{O}(\theta)$ for every trial, so that the state of all $N$ object samples is $\bm{\rho}_{O}(\theta)=\rho_{O}(\theta)^{\otimes N}$. In the task of localizing point sources, considered here, $\theta$ includes the spatial coordinates $\bar{r}_{k}=(\bar{x}_{k},\bar{y}_{k})$ of each source on the object plane. In this task, the entire fluorescence session time ($\sim 1$ ms) is split into time bins of short duration ($\sim 1$ ns) so that not more than one photon is detected in each time bin. The empty time bins are discarded, while the time bins with detections are numbered with $n=1,...,N$, where $N$ is the total number of photons detected. The point $r_{n}$ on the image plane where a photon was detected in the $n$th bin is considered as a result of a measurement of the $n$th sample of the field. That is, it is implied that the sources’ coordinates do not vary with time and the subsequent acts of spontaneous emission are statistically independent. Note that the exact times of photon detections may not be resolved by the camera of the microscope and are not necessary for the statistical model. Even if the probability of photon detection in one time bin is as small as $0.01$, the total number of detected photons can be rather high, typically $N\sim 10^{4}$. Specifically, $\rho_{O}(\theta)$ can be understood as the density operator of the electronic excitation in a time bin where a photon was detected. Alternatively, it can be understood as the density operator of a spontaneously emitted photon on the object plane.
@@
In the first stage, the probe is prepared in a state $\bm{\rho}_{P}$, which is, in the simplest case of independent identical systems, just a direct product of the states of each sample of the probe $\rho_{P}$: $\bm{\rho}_{P}=\rho_{P}^{\otimes N}$. However, more complicated entangled probes can be used for reaching the quantum-enhanced sensitivity [23], which we do not consider here. In optical microscopy, the probe is represented by the part of the optical field captured by the aperture of the microscope. For the spontaneous emission of fluorophores, its initial state is vacuum. Thus, its preparation simply means screening the field of view of the microscope from the background light in the frequency band used for imaging.
@@
In the second stage, the probe and object undergo an interaction, which is described in the most general form by a completely positive trace-preserving linear map $\mathcal{E}$, so that the state of all $N$ samples of the object and probe after the interaction is $\bm{\rho}_{OP}(\theta)=\mathcal{E}\left[\bm{\rho}_{O}(\theta)\otimes\bm{\rho}_{P}\right]$. This map can be made unitary on an extended state space by a Stinespring dilation, that is, by adding up an environment system with the initial state $\bm{\rho}_{E}$ and tracing out the environment after the interaction: $\bm{\rho}_{OP}(\theta)=\operatorname{Tr}_{E}\left\{\bm{U}\bm{\rho}_{O}(\theta)\otimes\bm{\rho}_{P}\otimes\bm{\rho}_{E}\bm{U}^{\dagger}\right\}$, where $\bm{U}$ is the unitary operator describing the interaction of the object, probe, and environment and $\dagger$ stands for Hermit conjugation. In the simplest case of independent identical interactions, both the state of the environment and the interaction operator are simple direct products of one-sample operators $\rho_{E}$ and $U_{\theta}$: $\bm{\rho}_{E}=\rho_{E}^{\otimes N}$ and $\bm{U}=U^{\otimes N}$. In optical microscopy of fluorophores, the environment is represented by the part of the optical field not captured by the aperture of the microscope and the degrees of freedom of the fluorophore other than the optical transition responsible for the spontaneous emission. The interaction consists in optical excitation of a transition belonging to the environment, from where the excitation is transmitted to the object, which spontaneously emits a photon, belonging to the probe or to the environment.
@@
Summing up the above relations, we write the conditional probability of the measurement outcome $\bm{r}$ under the condition that the parameter is $\theta$ as
$\bm{p}(\bm{r}|\theta)=\operatorname{Tr}\left\{\bm{\Lambda}(\bm{r})\bm{U}\bm{\rho}_{O}(\theta)\otimes\bm{\rho}_{P}\otimes\bm{\rho}_{E}\bm{U}^{\dagger}\right\},$ (2)
where the trace is taken over all $N$ samples of the three systems: object, probe, and environment. The value of the parameter $\theta$ is inferred from the random outcome $\bm{r}$ by a function $\hat{\theta}(\bm{r})$ known as the estimator, which is again a random variable. The estimator is called unbiased if its mean coincides with the true value of the parameter, $\langle\hat{\theta}(\bm{r})\rangle=\theta$, where the angular brackets denote the averaging with the conditional probability, Eq. (2). For any unbiased estimator, its covariance matrix $\operatorname{Cov}(\hat{\theta})_{jk}=\langle(\hat{\theta}_{j}-\theta_{j})(\hat{\theta}_{k}-\theta_{k})\rangle$ is lower bounded by the Cramér-Rao bound discussed in the Introduction for a single parameter and having a matrix form in the multiparameter case [18, 19]:
$\operatorname{Cov}(\hat{\theta})\geq\mathbf{F}^{-1}(\theta),$ (3)
where the matrix inequality $\mathrm{A}\geq\mathrm{B}$ for two Hermitian matrices $\mathrm{A}$ and $\mathrm{B}$ means that the matrix $\mathrm{A}-\mathrm{B}$ is positive semidefinite, i.e., all its eigenvalues are nonnegative [35], and $\mathbf{F}(\theta)$ is the Fisher information matrix, whose elements are defined as
[…]
Here and in the following, we assume that $\bm{p}(\bm{r}|\theta)$ is nonzero everywhere, which is true for an imaging system whose point-spread function is modeled by a Gaussian. Note that we use bold Roman font for matrices and bold italic font for scalars and vectors related to the entire set of $N$ samples.
In optical microscopy, the typical estimator for the location of a single emitter is the result of fitting the distribution of the field intensity on its image spot by a two-dimensional Gaussian function [9, 12]. In the case of multiple emitters, the Cramér-Rao bound is saturated by finding the parameter $\hat{\theta}(\bm{r})$ which maximizes $\bm{p}(\bm{r}|\theta)$ considered as a function of $\theta$ and known in this aspect as likelihood, which gives the name of “maximal likelihood” to the estimator [18, 19].
II.2 Localization efficiency
The diagonal elements of the matrix $\operatorname{Cov}(\hat{\theta})$ give variances of the components of $\theta$, $\operatorname{Cov}(\hat{\theta})_{ii}=\operatorname{Var}(\hat{\theta}_{i})$ and traditionally the entire diagonal of the covariance matrix is considered as the indicator of success in determining the unknown parameters. However, when tens or hundreds of sources are available, one might be interested in one number characterizing the efficiency of mass localization. We prefer to express this “localization efficiency” through the inverse of the parameter variance known as precision [36, 28], so that a high efficiency corresponds to a low position variance. This localization efficiency can be introduced in several alternative ways, depending on the order in which the averaging and inversion are performed.
@@ II.2 Localization efficiency
$H_{\textnormal{tot}}(\hat{\theta})=\frac{M}{\operatorname{Tr}\{\operatorname{Cov}(\hat{\theta})\}}.$ (5)
@@
$H_{\textnormal{tot}}(\hat{\theta})\leq\frac{M}{\operatorname{Tr}\{\mathbf{F}^{-1}(\theta)\}},$ (6)
@@
$\bar{F}_{\text{ind}}(\theta)=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{\mathbf{F}^{-1}(\theta)_{ii}}$ (8)
@@
Both $H_{\text{ind}}(\hat{\theta})$ and $\bar{F}_{\text{ind}}(\theta)$ share the drawback of being non-invariant with respect to rotations in the parameter space. Aiming at an invariant measure, we introduce one more measure of localization efficiency. First, we define the precision matrix [36] $\operatorname{Prec}(\hat{\theta})=\operatorname{Cov}(\hat{\theta})^{-1}$, whose diagonal entry $\operatorname{Prec}(\hat{\theta})_{ii}$ is larger than or equal to the individual precision $1/\operatorname{Cov}(\hat{\theta})_{ii}$. If we consider an orthogonal rotation in the parameter space, $\varphi=O^{T}\theta$, where $\varphi=\{\varphi_{1},...,\varphi_{M}\}$, both parameter vectors are understood as column vectors, $T$ denotes a transposition, and $O$ is an orthogonal matrix such that $\operatorname{Cov}(\hat{\varphi})=O^{T}\operatorname{Cov}(\hat{\theta})O$ is diagonal with $\hat{\varphi}=O^{T}\hat{\theta}$ an unbiased estimator for $\varphi$, then $\operatorname{Prec}(\hat{\varphi})_{ii}=1/\operatorname{Cov}(\hat{\varphi})_{ii}$. We will call $\varphi$ the eigenparameter of the localization task. Now, we define the third measure of localization efficiency as follows.
@@
$H_{\textnormal{eig}}(\hat{\theta})=\frac{1}{M}\operatorname{Tr}\{\operatorname{Prec}(\hat{\theta})\}.$ (9)
The upper bound for this quantity can be found from Eq. (3). For any two invertible Hermitian matrices $A$ and $B$, $A\geq B$ if and only if $A^{-1}\leq B^{-1}$ [35]. Note that the covariance, precision and Fisher information matrices are real symmetric by definition, and as such are Hermitian. Therefore, Eq. (3) can be rewritten as $\operatorname{Prec}(\hat{\theta})\leq\mathbf{F}(\theta)$. Taking a trace of both sides of this inequality, we obtain $H_{\text{eig}}(\hat{\theta})\leq\bar{F}(\theta)$, where
$\bar{F}(\theta)=\frac{1}{M}\operatorname{Tr}\left\{\mathbf{F}(\theta)\right\}$ (10)
is the average Fisher information. This bound is asymptotically saturated by the maximum-likelihood estimator. The matrix $O$ diagonalizing the covariance matrix coincides asymptotically with that diagonalizing the Fisher information matrix. However, the matrix $O$ is itself dependent on the unknown parameters and is typically unknown when the localization procedure starts.
@@
It should also be understood that the measures given by Definitions 2 and 3 are not intended for use in the fitting algorithms. The quantities $H_{\text{ind}}(\hat{\theta})$ and $H_{\text{eig}}(\hat{\theta})$ may be very large in the case where only one source is localized with very high precision but the positions of all other sources are totally unknown, which is not the best localization outcome. We assume that the fitting is realized by the maximum likelihood estimation and the measure of localization efficiency is applied afterwards to the results of such estimation. The evaluation of efficiency here is similar to that in quantum teleportation [37, 38]: the verifier knowing the exact positions of the sources gives a sample to the microscopist, who returns the estimated source positions, for which the verifier calculates the measure of efficiency, similar in this respect to the teleportation fidelity.
@@
$H_{\text{tot}}(\hat{\theta})\leq H_{\text{ind}}(\hat{\theta})\leq H_{\text{eig}}(\hat{\theta}).$ (11)
@@
Substituting $x_{i}=\operatorname{Cov}(\hat{\theta})_{ii}$, we obtain the first inequality. To prove the second inequality, consider an invertible real symmetric $M\times M$ matrix $A$ and a unit $M\times 1$ vector $v$. Let $O$ be an orthogonal matrix that brings $A$ to the diagonal form, $A=ODO^{T}$, where $D$ is diagonal. Then $v^{T}Av=\tilde{v}^{T}D\tilde{v}$, where $\tilde{v}=O^{T}v$. Since $\tilde{v}$ is also a unit vector, the numbers $\tilde{v}_{i}^{2}$ sum up to unity and can be considered as statistical weights. From the convexity of the function $f(x)=1/x$ on the interval $x\in(0,+\infty)$, as above, we obtain
$f\left(\sum\limits_{i=1}^{M}\tilde{v}_{i}^{2}D_{i}\right)\leq\sum\limits_{i=1}^{M}\tilde{v}_{i}^{2}f(D_{i}),$ (13)
or $(\tilde{v}^{T}D\tilde{v})^{-1}\leq\tilde{v}^{T}D^{-1}\tilde{v}$ and, as a consequence, $(v^{T}Av)^{-1}\leq v^{T}A^{-1}v$. Choosing a vector $v_{i}=\delta_{im}$ for some $1\leq m\leq M$, we obtain the inequality $1/A_{mm}\leq(A^{-1})_{mm}$. Substituting $A_{mm}=\operatorname{Cov}(\hat{\theta})_{mm}$ and summing up over $m$, we obtain the second inequality of Eq. (11).

### v3

@@ I Introduction
$\left(\Delta x\right)^{2}\geq 1/\underline{F},$ (1)
where $\underline{F}$ is the Fisher information on this parameter contained in the measurement data. Here and below underlined letters denote quantities related to the entire measurement session typically consisting of many photodetections. In the case of localization of a single emitter, the subsequent photodetections are statistically independent, and the total Fisher information carried by $N$ photons is simply $\underline{F}=NF_{\text{1phot}}$, where $F_{\text{1phot}}$ is the Fisher information on the emitter position carried by one photon. The Cramér-Rao bound can be asymptotically reached in the high $N$ limit when the estimator is given by the maximum likelihood estimation [18, 19]. Thus, the $1/\sqrt{N}$ scaling of the position standard deviation is a mere consequence of the additivity of Fisher information with respect to independent observations. The single-photon Fisher information can be calculated for various models of the PSF and the accompanying noise [20, 21] providing the Cramér-Rao bound for any given $N$. The variance of the estimated emitter position was shown to approach the Cramér-Rao bound for $N=10^{2}-10^{4}$ photons allowing thus for superresolution.
The same argument can be applied to determining the separation of two closely spaced emitters: the total Fisher information carried by $N$ photons on the separation is $\underline{F}=NF_{\text{sep,1phot}}$, where $F_{\text{sep,1phot}}$ is the information on the separation carried by one photon. This information tends to zero together with the separation of the emitters [22], however, for whatever small $F_{\text{sep,1phot}}$, one can make the total information arbitrarily large by collecting sufficiently large number of photons and thus determine the separation with arbitrary precision. In practice, a fluorophore emits a limited number of photons before photobleaching, and, therefore, individual localization of blinking emitters gives much better results than a simultaneous localization of multiple cofluorescent emitters within the width of PSF. The precision gain attained by making the sources blink is the key feature of SMLM. A question arises: Is it a general law of Nature that blinking emitters send more information on their locations than cofluorescent ones?
@@
The paper is structured as follows. In Sec. II we recast the formalism of quantum metrology [23, 24, 25, 26] in the context of localization of dim incoherent sources of light [22, 27, 28, 29, 30], extending a more traditional approach of classical metrology typically used in this context [20, 21, 31, 32, 33, 34], and introduce three measures of localization efficiency. In Sec. III, we prove the fundamental properties of the Fisher information matrix mentioned above. In Sec. IV, we compare two scenarios of cofluorescent and blinking sources and show that the latter always has an informational advantage. We illustrate the developed theory by an example of two incoherent sources in Sec. V and extend the results to the quantum Fisher information matrix in Sec. VI. The obtained results are summarized in the Conclusion.
@@ II.1 Cramér-Rao bound
The objective of quantum metrology is inference of an unknown parameter $\theta$, which may be a vector $\theta=\{\theta_{1},...,\theta_{M}\}$, encoded into the state of the object $\rho_{O}(\theta)$, which is a density operator in the Hilbert space of the object. For this purpose a generalized quantum measurement of the object is repeated $N$ times, the object being prepared in the same state $\rho_{O}(\theta)$ for every trial, so that the state of all $N$ object samples is $\underline{\rho}_{O}(\theta)=\rho_{O}(\theta)^{\otimes N}$. In the task of localizing point sources, considered here, $\theta$ includes the spatial coordinates $\bar{r}_{k}=(\bar{x}_{k},\bar{y}_{k})$ of each source on the object plane. In this task, the entire fluorescence session time ($\sim 1$ ms) is split into time bins of short duration ($\sim 1$ ns) so that not more than one photon is detected in each time bin. The empty time bins are discarded, while the time bins with detections are numbered with n=1,,Nn=1,...,N, where $N$ is the total number of photons detected. The point $r_{n}$ on the image plane where a photon was detected in the $n$th bin is considered as a result of a measurement of the $n$th sample of the field. That is, it is implied that the sources’ coordinates do not vary with time and the subsequent acts of spontaneous emission are statistically independent. Note that the exact times of photon detections may not be resolved by the camera of the microscope and are not necessary for the statistical model. Even if the probability of photon detection in one time bin is as small as $0.01$, the total number of detected photons can be rather high, typically $N\sim 10^{4}$. Specifically, $\rho_{O}(\theta)$ can be understood as the density operator of the electronic excitation in a time bin where a photon was detected. Alternatively, it can be understood as the density operator of a spontaneously emitted photon on the object plane.
Figure 1: Scheme for measuring the positions of two incoherent sources by imaging them in a microscope equipped with a single-photon avalanche diode (SPAD) array and estimating the positions from the measurement data. The dotted red lines show the rays that form the geometrical images at points $x_{1}$ and $x_{2}$, while the solid red line shows the field intensity distribution in the image plane as a sum of two point-spread functions centered at $x_{1}$ and $x_{2}$.
@@
In the first stage, the probe is prepared in a state $\underline{\rho}_{P}$, which is, in the simplest case of independent identical systems, just a direct product of the states of each sample of the probe $\rho_{P}$: $\underline{\rho}_{P}=\rho_{P}^{\otimes N}$. However, more complicated entangled probes can be used for reaching the quantum-enhanced sensitivity [23], which we do not consider here. In optical microscopy, the probe is represented by the part of the optical field captured by the aperture of the microscope, see Fig. 1. For the spontaneous emission of fluorophores, its initial state is vacuum. Thus, its preparation simply means screening the field of view of the microscope from the background light in the frequency band used for imaging.
@@
In the second stage, the probe and object undergo an interaction, which is described in the most general form by a completely positive trace-preserving linear map $\mathcal{E}$, so that the state of all $N$ samples of the object and probe after the interaction is $\underline{\rho}_{OP}(\theta)=\mathcal{E}\left[\underline{\rho}_{O}(\theta)\otimes\underline{\rho}_{P}\right]$. This map can be made unitary on an extended state space by a Stinespring dilation, that is, by adding up an environment system with the initial state $\underline{\rho}_{E}$ and tracing out the environment after the interaction: $\underline{\rho}_{OP}(\theta)=\operatorname{Tr}_{E}\left\{\underline{U}\underline{\rho}_{O}(\theta)\otimes\underline{\rho}_{P}\otimes\underline{\rho}_{E}\underline{U}^{\dagger}\right\}$, where $\underline{U}$ is the unitary operator describing the interaction of the object, probe, and environment and $\dagger$ stands for Hermit conjugation. In the simplest case of independent identical interactions, both the state of the environment and the interaction operator are simple direct products of one-sample operators $\rho_{E}$ and $U_{\theta}$: $\underline{\rho}_{E}=\rho_{E}^{\otimes N}$ and $\underline{U}=U^{\otimes N}$. In optical microscopy of fluorophores, the environment is represented by the part of the optical field not captured by the aperture of the microscope and the degrees of freedom of the fluorophore other than the optical transition responsible for the spontaneous emission. The interaction consists in optical excitation of a transition belonging to the environment, from where the excitation is transmitted to the object, which spontaneously emits a photon, belonging to the probe or to the environment.
In the third stage, the probe is measured giving the outcome $\underline{r}$, which is a combination of measurement outcomes $r_{n}$ for every sample of the probe: $\underline{r}=\{r_{1},...,r_{N}\}$. The probability of this outcome is given by the Born rule $\underline{p}(\underline{r}|\theta)=\operatorname{Tr}_{P}\left\{\underline{\Lambda}(\underline{r})\tilde{\underline{\rho}}_{P}(\theta)\right\}$, where $\tilde{\underline{\rho}}_{P}(\theta)=\operatorname{Tr}_{O}\left\{\underline{\rho}_{OP}(\theta)\right\}$ is the state of the probe with the object traced out, and $\underline{\Lambda}(\underline{r})$ is the positive-operator-valued measure (POVM) for the outcome $\underline{r}$. As before, for independent identical measurements, the POVM is a direct product of POVMs for every sample: $\underline{\Lambda}(\underline{r})=\otimes_{n=1}^{N}\Lambda(r_{n})$. Collective measurements of probe samples are studied in some scenarios of quantum metrology [23], but are not considered here. When the magnification of the microscope is sufficiently high, its field of view is covered by a large enough number of pixels, so that the measurement outcome can be represented by the continuous coordinate $r_{n}=\{x_{n},y_{n}\}$ of the detected photon. The effects of detector pixelation are studied in the literature [20], but are not considered here. The corresponding single-sample POVM is the projection on the single-photon state at point $r_{n}$: $\Lambda(r_{n})=|r_{n}\rangle\langle r_{n}|$, where $|r_{n}\rangle=a^{\dagger}(r_{n})|\text{vac}\rangle$, $a^{\dagger}(r_{n})$ being the photon creation operator at point $r_{n}$ and $|\text{vac}\rangle$ is the vacuum state of the field.
Summing up the above relations, we write the conditional probability of the measurement outcome $\underline{r}$ under the condition that the parameter is $\theta$ as
$\underline{p}(\underline{r}|\theta)=\operatorname{Tr}\left\{\underline{\Lambda}(\underline{r})\underline{U}\underline{\rho}_{O}(\theta)\otimes\underline{\rho}_{P}\otimes\underline{\rho}_{E}\underline{U}^{\dagger}\right\},$ (2)
where the trace is taken over all $N$ samples of the three systems: object, probe, and environment. The value of the parameter $\theta$ is inferred from the random outcome $\underline{r}$ by a function $\hat{\theta}(\underline{r})$ known as the estimator, which is again a random variable. The estimator is called unbiased if its mean coincides with the true value of the parameter, $\langle\hat{\theta}(\underline{r})\rangle=\theta$, where the angular brackets denote the averaging with the conditional probability, Eq. (2). For any unbiased estimator, its covariance matrix $\operatorname{Cov}(\hat{\theta})_{jk}=\langle(\hat{\theta}_{j}-\theta_{j})(\hat{\theta}_{k}-\theta_{k})\rangle$ is lower bounded by the Cramér-Rao bound discussed in the Introduction for a single parameter and having a matrix form in the multiparameter case [18, 19]:
$\operatorname{\mathbf{Cov}}(\hat{\theta})\geq\underline{\mathbf{F}}^{-1}(\theta),$ (3)
where the matrix inequality $\mathbf{A}\geq\mathbf{B}$ for two Hermitian matrices $\mathbf{A}$ and $\mathbf{B}$ means that the matrix $\mathbf{A}-\mathbf{B}$ is positive semidefinite, i.e., all its eigenvalues are nonnegative [35], and $\underline{\mathbf{F}}(\theta)$ is the Fisher information matrix, whose elements are defined as
$\underline{F}_{ij}(\theta)=\int\frac{1}{\underline{p}(\underline{r}|\theta)}\frac{\partial\underline{p}(\underline{r}|\theta)}{\partial\theta_{i}}\frac{\partial\underline{p}(\underline{r}|\theta)}{\partial\theta_{j}}d\underline{r}.$ (4)
Here and in the following, we assume that $\underline{p}(\underline{r}|\theta)$ is nonzero everywhere, which is true for an imaging system whose point-spread function is modeled by a Gaussian. Note that we use bold font for matrices, italic for their elements, Roman for the elements of function matrices, and underline, as above, for the quantities related to the entire set of $N$ samples.
In optical microscopy, the typical estimator for the location of a single emitter is the result of fitting the distribution of the field intensity on its image spot by a two-dimensional Gaussian function [9, 12]. In the case of multiple emitters, the Cramér-Rao bound is saturated by finding the parameter $\hat{\theta}(\underline{r})$ which maximizes $\underline{p}(\underline{r}|\theta)$ considered as a function of $\theta$ and known in this aspect as likelihood, which gives the name of “maximal likelihood” to the estimator [18, 19].
II.2 Localization efficiency
@@
$H_{\textnormal{tot}}(\hat{\theta})=\frac{M}{\operatorname{Tr}\{\operatorname{\mathbf{Cov}}(\hat{\theta})\}}$. (5)
@@
$H_{\textnormal{tot}}(\hat{\theta})\leq\frac{M}{\operatorname{Tr}\{\underline{\mathbf{F}}^{-1}(\theta)\}},$ (6)
@@
$\bar{F}_{\text{ind}}(\theta)=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{\left[\underline{\mathbf{F}}^{-1}(\theta)\right]_{ii}}$ (8)
@@
[…]
@@
[…]
The upper bound for this quantity can be found from Eq. (3). For any two invertible Hermitian matrices $\mathbf{A}$ and $\mathbf{B}$, $\mathbf{A}\geq\mathbf{B}$ if and only if $\mathbf{A}^{-1}\leq\mathbf{B}^{-1}$ [35]. Note that the covariance, precision and Fisher information matrices are real symmetric by definition, and as such are Hermitian. Therefore, Eq. (3) can be rewritten as $\operatorname{\mathbf{Prec}}(\hat{\theta})\leq\underline{\mathbf{F}}(\theta)$. Taking a trace of both sides of this inequality, we obtain $H_{\text{eig}}(\hat{\theta})\leq\bar{F}(\theta)$, where
$\bar{F}(\theta)=\frac{1}{M}\operatorname{Tr}\left\{\underline{\mathbf{F}}(\theta)\right\}$ (10)
is the average Fisher information. This bound is asymptotically saturated by the maximum-likelihood estimator. The matrix $\mathbf{O}$ diagonalizing the covariance matrix coincides asymptotically with that diagonalizing the Fisher information matrix. However, the matrix $\mathbf{O}$ is itself dependent on the unknown parameters and is typically unknown when the localization procedure starts.
@@
[…]
@@
[…]
@@
Substituting $x_{i}=\operatorname{\mathbf{Cov}}(\hat{\theta})_{ii}$, we obtain the first inequality. To prove the second inequality, consider an invertible real symmetric $M\times M$ matrix $\mathbf{A}$ and a unit $M\times 1$ vector $v$. Let $\mathbf{O}$ be an orthogonal matrix that brings $\mathbf{A}$ to the diagonal form, $\mathbf{A}=\mathbf{O}\mathbf{D}\mathbf{O}^{T}$, where $\mathbf{D}$ is diagonal. Then $v^{T}\mathbf{A}v=\tilde{v}^{T}\mathbf{D}\tilde{v}$, where $\tilde{v}=\mathbf{O}^{T}v$. Since $\tilde{v}$ is also a unit vector, the numbers $\tilde{v}_{i}^{2}$ sum up to unity and can be considered as statistical weights. From the convexity of the function $f(x)=1/x$ on the interval $x\in(0,+\infty)$, as above, we obtain
$f\left(\sum\limits_{i=1}^{M}\tilde{v}_{i}^{2}D_{ii}\right)\leq\sum\limits_{i=1}^{M}\tilde{v}_{i}^{2}f(D_{ii}),$ (13)
or $(\tilde{v}^{T}\mathbf{D}\tilde{v})^{-1}\leq\tilde{v}^{T}\mathbf{D}^{-1}\tilde{v}$ and, as a consequence, $(v^{T}\mathbf{A}v)^{-1}\leq v^{T}\mathbf{A}^{-1}v$. Choosing a vector $v_{i}=\delta_{im}$ for some $1\leq m\leq M$, we obtain the inequality $1/A_{mm}\leq(\mathbf{A}^{-1})_{mm}$. Substituting $A_{mm}=\operatorname{Cov}(\hat{\theta})_{mm}$ and summing up over $m$, we obtain the second inequality of Eq. (11).